Fermi National Laboratory


Accelerator Update for March 11, 2005 - March 14, 2005

Friday March 11
The day shift began with experts investigating the Tevatron (TeV) Quench Protection Monitor (QPM) system, and with MiniBooNE, NuMI, and the Meson MCenter experiments taking beam.

Operations established beam to the Meson MTest experiment at 9:45 AM.

An expert reported at 1:06 PM that the QPM system was fixed.

Operators put the TeV through a dry squeeze at 2:25 PM.

Operations reported at 3:46 PM that Preaccelerator experts had turned on the I- Source and that voltage glitches on the I- Source had caused H- Source trips. Experts made adjustments.

At 5:04 PM, Operations reported that the NuMI LCW resistivity had been going bad. LCW experts came in just before shift change and valved in two deionizing bottles. NuMI will remain off until the resistivity gets better.

Operations put the TeV into shot setup at 5:12 PM. Shot setup disables beam to NuMI and Switchyard (SY).

Operations established store 4041 at 9:32 PM with an initial luminosity of 84.9E30. Pbar resumed stacking.

MCenter and MTest resumed taking beam 7:41 PM.

Saturday March 12
The midnight shift began with Operations monitoring store 4041, with operators preparing the Recycler to stash antiprotons, and with beam to MiniBooNE, MCenter, and MTest.

Operators put the Recycler into shot setup at 1:20 AM, but problems in the Main Injector (MI) sequencer lost the last transfer of antiprotons.

Pbar resumed stacking at 2:18 AM.

MCenter resumed taking beam at 2:36 AM.

MTest experimenters resumed taking beam at 8:13 AM.

The LCW experts reported at 9:07 AM that the resistivity was good enough for the NuMI power supplies to be turned back on and for NuMI to resume running. At 9:21 AM, the NuMI experts agreed.

NuMI experts began a study period at 9:50 AM.

At 11:36 AM, MI began suffering from beam aborts. An expert came in and fixed a timing problem.

At 7:51 PM, a NuMI power supply (HV101A) tripped off and wouldn't reset. Operators called in an expert.

Operations terminated store 4041 at 8:01 PM.

Operations put the TeV into shot setup at 8:14 PM.

Operations established store 4042 at 10:46 PM with an initial luminosity of 102.29E30. Pbar resumed stacking.

MCenter resumed taking beam at 11:38 PM.

At 11:49 PM, Operations reported a vacuum problem in the P3 transfer line to SY. Operators found an open vacuum valve. This will hold off beam to SY until the vacuum gets better.

Sunday March 13
The midnight shift began with Operations monitoring store 4042, with Pbar stacking, with MiniBooNE taking beam, but with SY off due to vacuum problems.

Operations reported at 1:31 AM that the vacuum had improved enough in the P3 line to send beam to MCenter.

At 1:39 AM, a TeV sector B1 air pressure came into alarm. Operators responded and found a broken fitting on an air compressor. They made a temporary repair, which saved the store and held until the Duty Mechanic arrived and made a permanent fix.

At 1:52 AM, a Booster RF station (BRF5) tripped off due to an LCW problem. An operator reset it, but it tripped off again. Operators investigated and found LCW spraying out. They valved it off and bypassed the station.

At 3:20 AM, Operations reported that beam was not making it into the Accumulator. Operators investigated and discovered that a kicker (A:IKIK) power supply wasn't firing. Operators replaced a trigger module for the supply at 3:47 AM and stacking resumed.

Operations put the Recycler into shot setup and began stashing antiprotons at 8:39 AM.

Stashing ended and Pbar resumed stacking at 9:35 AM.

At 9:50 AM, A:IKIK stopped working again. An operator called in and expert. By 12:53 PM, he had replaced a bad cable and stacking resumed.

At 4:11 PM, operators and cryo system technicians investigated a failing wet engine for TeV sector B4. They found a problem with the flywheel. Cryo techs made some adjustments that they hoped would keep the engine running until the end of the store.

Monday March 14
The midnight shift began with Operations monitoring store 4042 and the Recycler stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

NuMI experts reported at 12:20 AM that they were done with the beam for the night.

Operations terminated store 4042 at 6:22 AM.

Cryo techs began working on the TeV B4 wet engine at 7:02 AM. They finished their work and Operations put the TeV into shot setup at 7:44 AM.

The Plan for Monday
The plan for today is to establish a new store and then resume stacking and stashing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 3/14/2005   email Fermilab